26 Feb (Thu)

  • LHCb has migrated to a new catalog architecture on Monday (Dirac File Catalog). The migration went all successful. Many thanks to IT/DB and IT/PES colleagues for their support.
  • Distributed computing dominated by Monte Carlo and user activities.
  • T0:
    • Q: VOMRS move for LHCb date? No reply to some questions yet.
  • T1:
    • Multiple T1 (and T2D) sites in scheduled DT on Tuesday with their storage. LHCb would like to discuss a possibility to align such DTs better in sequence.

23 Feb (Mo)

  • LHCb dist computing activities drained b/c of migration to new file catalog system. We expect to come back from this DT this afternoon.
  • T0:
    • Migration of DBOD DBs together with the above migration (other than FC)
  • T1:

19 February 2015 (Thursday)

  • User jobs and MC at the moment.
  • T0: NTR
  • T1: NTR

16 February 2015 (Monday)

  • Still finalising last few file statuses of Stripping21 campaign. User jobs and MC at the moment.
  • T0: NTR
  • T1: NTR

12 February 2015 (Thursday)

  • Still finalising last few file statuses of Stripping21 campaign. User jobs and MC at the moment. Expect MC to ramp up soon.
  • T0: CERN job failures ticket still open (GGUS:111565) and on us to understand where the job time limit problem is.
  • T1: Thank you to T1s for bearing with us this week while we killed jobs from a user who had submitted a large number of tape requests.

9 February 2015 (Monday)

  • "Legacy Run1 Stripping" still looking at recovery from lost files. Large MC campaign upcoming + User jobs but not much happening right now.
  • T0: Many of the job failures at CERN appear to be related to batch queue timeleft mechanism. Investigating still. (GGUS:https://ggus.eu/?mode=ticket_info&ticket_id=111565)
  • T1: NTR

5 February 2015 (Thursday)

  • "Legacy Run1 Stripping" waiting on FTS tranfers and recovery from lost files. Large MC campaign upcoming + User jobs but not much happening right now.
  • T0: Lot of jobs failing at CERN (GGUS:https://ggus.eu/?mode=ticket_info&ticket_id=111565)
  • T1: SARA : any news about the INCIDENT report ?.

2 February 2015 (Monday)

  • "Legacy Run1 Stripping" waiting on FTS tranfers and recovery from lost files. Large MC campaign upcoming + User jobs but not much happening right now.
  • T0: NTR
  • T1: Downtime at RAL and forthcoming intervention on IN2P3 on CE transparent for user.

-- JoelClosier - 2015-06-01

Edit | Attach | Watch | Print version | History: r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r1 - 2015-06-01 - JoelClosier
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LHCb All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback